OpenStudy (anonymous):

f : x -> 3 - x what is the value of f(-4)?

OpenStudy (anonymous):

That just means put in -4 wherever you see an x.

OpenStudy (anonymous):

i got the answer 7 is that correct?

OpenStudy (anonymous):

Yes. f(x) = 3-x so f(-4) = 3 - (-4) = 7

OpenStudy (anonymous):

thanks